# On the well-posedness of SPDEs with locally Lipschitz coefficients

### v2

@@ 3. Proof of Theorem 3
and the rate of convergence does not depend on $t\in(0\,,T)$ nor on $x\in\mathbb{R}$. The random field $u$ is predictable since it is a limit of predictable random fields. Therefore it remains to prove that, for every $t\in(0\,,T)$ and $x\in\mathbb{R}$,
$\lim_{N\to\infty}\mathcal{I}^{N}_{b}(t\,,x)=\mathcal{I}_{b}(t\,,x)\quad\text{and}\quad\lim_{N\to\infty}\mathcal{I}^{N}_{\sigma}(t\,,x)=\mathcal{I}_{\sigma}(t\,,x),$ (3.6)
@@
where both of the limits hold in $L^{2}(\Omega)$, and the random fields $\mathcal{I}_{b}$, $\mathcal{I}_{\sigma}$, $\mathcal{I}^{N}_{b}$, and $\mathcal{I}^{N}_{\sigma}$ were defined respectively in (1.4) and (2.5). Thanks to (3.5) and (2.4), this proves that $u$ is a mild solution to (1.1).
Observe that $|b_{N}(t\,,z)-b(t\,,z)|\leq\mathbf{1}_{\mathbb{R}\setminus[-\exp(N),\exp(N)]}(z)(|b(t\,,z)|+|b_{N}(t\,,z)|)\leq 2L_{b}(1+|z|)\mathbf{1}_{\mathbb{R}\setminus[-\exp(N),\exp(N)]}(z)$ for all $z\in\mathbb{R}$ and $t,N>0$. This yields
$\left\|\mathcal{I}^{N}_{b}(t\,,x)-\int_{(0,t)\times\mathbb{R}}p_{t-s}(y-x)b(s\,,u_{N}(s\,,y))\,\mathrm{d}s\,\mathrm{d}y\right\|_{2}$
@@
If $X\geq 0$ is a random variable and $A>0$ is a constant, then we apply the Cauchy-Schwarz inequality and Chebyshev’s inequality back to back in order to find that $\mathrm{E}(X^{2};X>A)\leq\sqrt{\mathrm{E}(X^{4})\mathrm{P}\{|X|>A\}}\leq A^{-2}\mathrm{E}(X^{4}),$ Therefore, Propositions 4 and 5 assure us that
$\lim_{N\to\infty}\adjustlimits{\sup}_{s\in(0,T)}{\sup}_{y\in\mathbb{R}}\mathrm{E}\left(1+|u_{N}(s\,,y)|^{2};|u_{N}(s\,,y)|>\mathrm{e}^{N}\right)=0,$
and hence $\mathcal{I}^{N}_{b}(t\,,x)-\int_{(0,t)\times\mathbb{R}}p_{t-s}(y-x)b(s\,,u_{N}(s\,,y))\,\mathrm{d}s\,\mathrm{d}y\to 0$ as $N\to\infty$, where the convergence takes place in $L^{2}(\Omega)$. Therefore, the first assertion of (3.6) would follow once we can show that
@@
where the convergence takes place in $L^{2}(\Omega)$. Since $b$ has at-most linear growth and $\|b(s\,,u_{N}(s\,,y))-b(s\,,u(s\,,y))\|_{2}\leq\|b(s\,,u_{N}(s\,,y))\|_{2}+\|b(s\,,u(s\,,y))\|_{2},$ Propositions 4 and 5 ensure that $\|b(s\,,u_{N}(s\,,y))-b(s\,,u(s\,,y))\|_{2}$ is bounded uniformly in $s\in(0\,,T)$, $N>0$, and $y\in\mathbb{R}$. Thanks to (3.5), uniform integrability, and the continuity of $b$,
@@
$\left\|\mathcal{I}^{N}_{\sigma}(t\,,x)-\int_{(0,t)\times\mathbb{R}}p_{t-s}(y-x)\sigma(s\,,u_{N}(s\,,y))\,W(\mathrm{d}s\,\mathrm{d}y)\right\|_{2}^{2}$
@@
As before, Propositions 4 and 5 assure us that
$\lim_{N\to\infty}\adjustlimits{\sup}_{s\in(0,T)}{\sup}_{y\in\mathbb{R}}\mathrm{E}\left(1+|u_{N}(s\,,y)|^{2};|u_{N}(s\,,y)|>\mathrm{e}^{N}\right)=0,$
and hence $\mathcal{I}^{N}_{\sigma}(t\,,x)-\int_{(0,t)\times\mathbb{R}}p_{t-s}(y-x)\sigma(s\,,u_{N}(s\,,y))\,W(\mathrm{d}s\,\mathrm{d}y)\to 0$ in $L^{2}(\Omega)$ as $N\to\infty$. Therefore, we are left to show that
@@
Propositions 4 and 5 ensure that $\|\sigma(s\,,u_{N}(s\,,y))-\sigma(s\,,u(s\,,y))\|_{2}$ is bounded uniformly in $s\in(0\,,T)$, $N>0$, and $y\in\mathbb{R}$. Thanks to (3.5), uniform integrability, and the continuity of $\sigma$, $\lim_{N\to\infty}\sigma(s\,,u_{N}(s\,,y))=\sigma(s\,,u(s\,,y))$ in $L^{2}(\Omega)$, for every $s>0$ and $y\in\mathbb{R}^{d}$. Therefore,
@@
which converges to 0 as $N\rightarrow\infty$ by dominated convergence, for all $t\in(0\,,T)$ and $x\in\mathbb{R}$. The triangle inequality now yields (3.7). This verifies the second assertion of (3.6) and completes the proof. ∎
@@
As in the proof of existence, we only give the details of the proof in the case that $L_{\sigma}>0$, since the remaining case is similar.
Let $u,v$ two solutions to (1.1) with same initial condition $u_{0}$ satisfying Assumption 1. Recall that $u$ and $v$ satisfy the mild formulation (1.3) with $\sigma$ and $b$ satisfying Assumption 2. Then, using Burkholder-Davis-Gundy inequality (see [minicourse]), we find that $\|u(t\,,x)-v(t\,,x)\|_{k}\leq I_{1}+I_{2}+I_{3}+I_{4}$ for all $t>0$, $x\in\mathbb{R}$, and $k\geq 1$, where
@@
$I_{2}^{2}$ $=8k\int_{0}^{t}\mathrm{d}s\int_{-\infty}^{\infty}\mathrm{d}y\ [p_{t-s}(y-x)]^{2}\left\|(\sigma(s\,,u(s\,,y))-\sigma(s\,,v(s\,,y))){\bf 1}_{A_{N}(s,y)}\right\|_{k}^{2},$
@@
Observe that the moment bounds obtained in Propositions 4 and 5 only use the linear growth constants $L_{\sigma}$ and $L_{b}$. Therefore, they also hold when $u_{N}$ is replaced by $u$ or $v$. The same is true for the tail estimates obtained Proposition 6, which also hold replacing $u_{N+1}$ by $u$ or $v$. Therefore, we can proceed as in the proof of existence but appeal to the local Lipschitz condition on $b$ and $\sigma$; that is, we recall (2.1) and write
@@
$\|u(t\,,x)-v(t\,,x)\|_{k}$ $\leq\mathrm{e}^{\beta t}\left[\frac{L_{N,b}}{\beta}+\frac{2\sqrt{k}L_{N,\sigma}}{\beta^{1/4}}\right]\mathcal{N}_{k,\beta,T}(u-v)$
$\quad+\frac{L_{b}C}{256L_{\sigma}^{4}k^{2}}\,\exp\left(512L_{\sigma}^{4}k^{2}t-\frac{N^{3/2}}{2k(512t)^{1/2}L_{\sigma}^{2}}\right)$
$\quad+\sqrt{32k}L_{\sigma}C\mathrm{e}^{512L_{\sigma}^{4}k^{2}t}\exp\left(-\frac{N^{3/2}}{2k(512t)^{1/2}L_{\sigma}^{2}}\right)t^{1/4},$
uniformly for all $\beta>0$, $t\in(0,T)$, $x\in\mathbb{R}$, and $N,k\geq c$, where $C=4(\|u_{0}\|_{L^{\infty}(\mathbb{R})}+1)$ and $c=c(\|u_{0}\|_{L^{\infty}(\mathbb{R})}\,,L_{\sigma}\,,T\,,L_{b})>0$. Then, choosing $\beta=512L_{N,\sigma}^{4}k^{6}$ and $k=\sqrt{L_{N,b}}$, and proceeding as in the proof of existence, we get that
$\adjustlimits{\sup}_{t\in(0,T)}{\sup}_{x\in\mathbb{R}}\|u(t\,,x)-v(t\,,x)\|_{2}\lesssim L_{N,b}^{-1}\exp\left(-\dfrac{N^{3/2}}{2(512T)^{1/2}L_{\sigma}^{2}\sqrt{L_{N,b}}}\right),$
uniformly for all $N\geq c^{\prime}=c^{\prime}(\|u_{0}\|_{L^{\infty}(\mathbb{R})}\,,L_{\sigma}\,,L_{b}\,,T)>0$. The liminf of the right-hand side of the latter expression converges to zero as $N$ tends to infinity [Lemma 7]. Thus, the proof of uniqueness is completed. ∎

### v3

@@ 2. Truncation, and preliminary results
and the rate of convergence does not depend on $t\in(0\,,T)$ nor on $x\in\mathbb{R}$. As a direct consequence $u$ is $L^{k}(\Omega)$-continuous. This ensures that $u$ has a predictable version. Therefore it remains to prove that, for every $t\in(0\,,T)$ and $x\in\mathbb{R}$,
$\lim_{N\to\infty}\mathcal{I}^{N}_{b_{N}}(t\,,x)=\mathcal{I}_{b}(t\,,x)\quad\text{and}\quad\lim_{N\to\infty}\mathcal{I}^{N}_{\sigma_{N}}(t\,,x)=\mathcal{I}_{\sigma}(t\,,x),$ (3.15)
@@
where both of the limits hold in $L^{2}(\Omega)$, and the random fields $\mathcal{I}_{b}$, $\mathcal{I}_{\sigma}$, $\mathcal{I}^{N}_{b}$, and $\mathcal{I}^{N}_{\sigma}$ were defined in (1.4) and (2.3). Thanks to (3.14) and (2.2), this proves that $u$ is a mild solution to (1.1).
Observe that
$\begin{split}|b_{N}(t\,,z)-b(t\,,z)|&\leq\mathbf{1}_{\mathbb{R}\setminus[-\exp(N),\exp(N)]}(z)(|b(t\,,z)|+|b_{N}(t\,,z)|)\\ &\leq 2L_{b}(1+|z|)\mathbf{1}_{\mathbb{R}\setminus[-\exp(N),\exp(N)]}(z),\end{split}$
for all $z\in\mathbb{R}$ and $t,N>0$. This yields
$\left\|\mathcal{I}^{N}_{b_{N}}(t\,,x)-\int_{(0,t)\times\mathbb{R}}p_{t-s}(y-x)b(s\,,u_{N}(s\,,y))\,\mathrm{d}s\,\mathrm{d}y\right\|_{2}$
@@
If $X\geq 0$ is a random variable and $A>0$ is a constant, then we apply the Cauchy-Schwarz inequality and Chebyshev’s inequality back to back in order to find that
$\mathrm{E}(X^{2};X>A)\leq\sqrt{\mathrm{E}(X^{4})\mathrm{P}\{|X|>A\}}\leq A^{-2}\mathrm{E}(X^{4}).$
Therefore, Propositions 2.1 and 2.3 assure us that
$\lim_{N\to\infty}\sup_{s\in(0,T)}\sup_{y\in\mathbb{R}}\mathrm{E}\left(1+|u_{N}(s\,,y)|^{2};|u_{N}(s\,,y)|>\mathrm{e}^{N}\right)=0,$
and hence $\mathcal{I}^{N}_{b_{N}}(t\,,x)-\int_{(0,t)\times\mathbb{R}}p_{t-s}(y-x)b(s\,,u_{N}(s\,,y))\,\mathrm{d}s\,\mathrm{d}y\to 0$ as $N\to\infty$, where the convergence takes place in $L^{2}(\Omega)$. Therefore, the first assertion of (3.15) would follow once we can show that
@@
where the convergence takes place in $L^{2}(\Omega)$. Since $b$ has at-most linear growth and $\|b(s\,,u_{N}(s\,,y))-b(s\,,u(s\,,y))\|_{2}\leq\|b(s\,,u_{N}(s\,,y))\|_{2}+\|b(s\,,u(s\,,y))\|_{2},$ Propositions 2.1 and 2.3 ensure that $\|b(s\,,u_{N}(s\,,y))-b(s\,,u(s\,,y))\|_{2}$ is bounded uniformly in $s\in(0\,,T)$, $N>0$, and $y\in\mathbb{R}$. Thanks to (3.14), uniform integrability, and the continuity of $b$,
@@
$\left\|\mathcal{I}^{N}_{\sigma_{N}}(t\,,x)-\int_{(0,t)\times\mathbb{R}}p_{t-s}(y-x)\sigma(s\,,u_{N}(s\,,y))\,W(\mathrm{d}s\,\mathrm{d}y)\right\|_{2}^{2}$
@@
As before, Propositions 2.1 and 2.3 assure us that
$\lim_{N\to\infty}\sup_{s\in(0,T)}\sup_{y\in\mathbb{R}}\mathrm{E}\left(1+|u_{N}(s\,,y)|^{2};|u_{N}(s\,,y)|>\mathrm{e}^{N}\right)=0,$
and hence $\mathcal{I}^{N}_{\sigma_{N}}(t\,,x)-\int_{(0,t)\times\mathbb{R}}p_{t-s}(y-x)\sigma(s\,,u_{N}(s\,,y))\,W(\mathrm{d}s\,\mathrm{d}y)\to 0$ in $L^{2}(\Omega)$ as $N\to\infty$. Therefore, we are left to show that
@@
Propositions 2.1 and 2.3 ensure that $\|\sigma(s\,,u_{N}(s\,,y))-\sigma(s\,,u(s\,,y))\|_{2}$ is bounded uniformly in $s\in(0\,,T)$, $N>0$, and $y\in\mathbb{R}$. Thanks to (3.14), uniform integrability, and the continuity of $\sigma$, $\lim_{N\to\infty}\sigma(s\,,u_{N}(s\,,y))=\sigma(s\,,u(s\,,y))$ in $L^{2}(\Omega)$, for every $s>0$ and $y\in\mathbb{R}^{d}$. Therefore,
@@
which converges to 0 as $N\rightarrow\infty$ by dominated convergence, for all $t\in(0\,,T]$ and $x\in\mathbb{R}$. The triangle inequality now yields (3.16). This verifies the second assertion of (3.15) and completes the proof of Theorem 1.4 in the case that $L_{\sigma}>0$. It remains to study the same problem when $\sigma$ is bounded.
The proof in the case that $\sigma$ is bounded is structurally similar to the derivation of the first part, except we replace the tail bound (2.10) with (2.11) in order to obtain the following variant of (3.13):
$\limsup_{N\to\infty}\mathrm{e}^{-2N}\log\adjustlimits{\sup}_{t\in(0,T]}{\sup}_{x\in\mathbb{R}}\|u_{N+1}(t\,,x)-u_{N}(t)\|_{k}<0\qquad\forall k\geq 1.$
The remainder of our derivation of the $\sigma$-bounded case is the same as in the first portion of the proof $L_{\sigma}>0$] and therefore omitted. This completes the proof. ∎
@@
We only give an outline of the proof in the case that $L_{\sigma}>0$, since the proof of uniqueness is essentially a simplied modification of the proof of existence, and because the other case where $\sigma$ is bounded is similar.
Let $u,v$ two solutions to (1.1) with same initial condition $u_{0}$ satisfying Assumption 1.1. Recall that $u$ and $v$ satisfy the mild formulation (1.3) with $\sigma$ and $b$ satisfying Assumption 1.2. Then, using Burkholder-Davis-Gundy inequality (see [minicourse]), we find that for all $t>0$, $x\in\mathbb{R}$, and $k\geq 1$,
$\|u(t\,,x)-v(t\,,x)\|_{k}\leq I_{1}+I_{2}+I_{3}+I_{4},$
where
@@
$I_{2}^{2}$ $=4k\int_{0}^{t}\mathrm{d}s\int_{-\infty}^{\infty}\mathrm{d}y\ [p_{t-s}(y-x)]^{2}\left\|(\sigma(s\,,u(s\,,y))-\sigma(s\,,v(s\,,y))){\bf 1}_{A_{N}(s,y)}\right\|_{k}^{2},$
@@
Observe that the moment bounds obtained in Propositions 2.1 and 2.3 only use the linear growth constants $L_{\sigma}$ and $L_{b}$. Therefore, they also hold when $u_{N}$ is replaced by $u$ or $v$. The same is true for the tail estimates obtained Proposition 2.4, which also hold replacing $u_{N+1}$ by $u$ or $v$. Therefore, we can proceed as in the proof of existence but appeal to the local Lipschitz condition on $b$ and $\sigma$; that is, we recall (1.5) and write
@@
$\|u(t\,,x)-v(t\,,x)\|_{k}$ $\leq\mathrm{e}^{\beta t}\left[L_{N,b}\beta^{-1}+2\sqrt{k}L_{N,\sigma}\beta^{-1/4}\right]\mathcal{N}_{k,\beta,T}(u-v)$
$+C\left[\frac{L_{b}}{128L_{\sigma}^{4}k^{2}}+\sqrt{128k}T^{1/4}L_{\sigma}\right]\exp\left(512L_{\sigma}^{4}k^{2}t-\frac{N^{3/2}}{128kL_{\sigma}^{2}\sqrt{t}}\right),$
uniformly for all $\beta>0$, $t\in(0,T)$, $x\in\mathbb{R}$, and $N,k\geq c$, where $C=4(\|u_{0}\|_{L^{\infty}(\mathbb{R})}+1)$ and $c=c(\|u_{0}\|_{L^{\infty}(\mathbb{R})}\,,L_{\sigma}\,,T\,,L_{b})>0$. Now we fix the parameters $\beta$ and $k$ as in (3.9) and analyze the preceding exactly as was done in the proof of existence in order to conclude that
$\sum_{N=1}^{\infty}\adjustlimits{\sup}_{t\in(0,T]}{\sup}_{x\in\mathbb{R}}\|u(t\,,x)-v(t\,,x)\|_{2}<\infty\qquad\forall T>0.$
Since the summand does not depend on $N$, it must be zero. This concludes the proof. ∎
Acknowledgements
We thanks Professors Raluca Balan and Tusheng Zhang for their encouraging remarks and for pointing out a serious error in an earlier draft. The correction of that error led to Assumption 1.3.